OpenStudy (anonymous):

sooooo confused Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)) a-40x + 24 b-40x + 31 c-40x + 64 d-40x + 70

OpenStudy (anonymous):

f(x) is a function of f, so it's assigning the value in the brackets to each "x" in it's equation. If we took f(2) we'd get -4(2) + 7. But, here it's f(g(x)), so we take g(x) and put it wherever we see an x in f(x). So f(g(x)) = -4(g(x)) + 7, since we know g(x) = 10x - 6, we can put that in the equation and neaten it up a little.

OpenStudy (anonymous):

-4(10x-6(x)) + 7 ? or would i -4(10x(x)) +7-6?

OpenStudy (anonymous):

Since \[g(x)=10x-6\] we can simply put in \[10x-6\] wherever we see an x in the f(x). So it would be; \[f(g(x)) = -4(10x-6) + 7\]

OpenStudy (anonymous):

-40x + 24 + 7=====-40x + 31

OpenStudy (anonymous):

i understand now thank you!

OpenStudy (anonymous):

Quite welcome!